OpenStudy (anonymous):

I am not sure I understand the following instructions on my book (Calculus: The Classic Edition, by Swokowski): "[a] Use Definition (3.1) to find the slope of the tangent line to the graph of the equation at the point of x-coordinate a. [b] Find an equation of the tangent line at P. [c] Sketch the graph and the tangent line at P." Then, I am given: y = square root of x; P(4, 2) The Definition (3.1) refers to the following: limit when h approaches 0 of [ f(a + h) - f(a) ] / h The answer book says that [a] is 1/(2*sqrt(a)) and [b] is y = (1/4)x + 1

OpenStudy (anonymous):

\[\lim_{h\to 0}\frac{\sqrt{x+h}-\sqrt{x}}{h}\] is a start

OpenStudy (anonymous):

I have figured out part [b] although I'm not understanding part [a].

OpenStudy (anonymous):